pooja195 (pooja195):

Margie is responsible for buying a week's supply of food and medication for the dogs and cats at a local shelter. The food and medication for each dog costs twice as much as those supplies for a cat. She needs to feed 164 cats and 24 dogs. Her budget is $4240. How much can Margie spend on each dog for food and medication? @iambatman i need help finding the equations i dont know how to do it without a verbal model

OpenStudy (anonymous):

Well you need to just state what your x and y values will be and then it's a breeze :P, can you do that?

pooja195 (pooja195):

x=dogs y-cats? right?

OpenStudy (anonymous):

Yes that works, x we'll say is dog food, y will be cat food

OpenStudy (compassionate):

164x + 2(24y) = 4240

OpenStudy (anonymous):

So we know 24x+164y = 4240 but we have another piece of information we can use, can you find it

OpenStudy (compassionate):

ITT: iambatman and Compassionate are trying to show Pooja how to solve the same equation dffernetly.

pooja195 (pooja195):

xD no wonder why this is confusing :P is the other equation d=2c?

OpenStudy (anonymous):

What are d and c

OpenStudy (compassionate):

dog and cat of course

pooja195 (pooja195):

xD i had a diffrent way of labeling it :P

OpenStudy (anonymous):

Oh ok, well lets keep the variables consistent, it's like when kids use different subscripts when using conservation of energy...uhg so annoying never works out :P

pooja195 (pooja195):

xD ok so x=2y?

OpenStudy (anonymous):

So you'll have x = 2y, yeah :P

OpenStudy (anonymous):

I'm sure you can finish it off from there

pooja195 (pooja195):

haha yep thanks :) both of u :)

OpenStudy (compassionate):

I think we're just conusing her Pooja, it's going to be a system of equations

OpenStudy (anonymous):

I think she knows, we just did a system of equations problem before this

OpenStudy (anonymous):

Or I hope she does xD

Nnesha (nnesha):

lO_Oks familiar pO_Oja jw

pooja195 (pooja195):

nnesha chup karo 164c + 24(2c) = 4240 164c + 48c = 4240 212c = 4240 c = 20 i did it again xD d n c d = 2c d = 2(20) = 40 is that right?

OpenStudy (anonymous):

c's and d's again! But yes it looks good :)

pooja195 (pooja195):

haha lol i cant help it thanks thought :D

pooja195 (pooja195):

*though